OpenStudy (anonymous):

Solve: if it is 68 degrees Fahrenheit what is it in celsius? Use this formula to help you c=(F-32)x5 divide 9 This is my 5th grade homework PLZZ HELP ANSWER

OpenStudy (anonymous):

C=(68-32)X5 just plug in your number for f which stands for fahrenheit. 68-32=36 36*5=180 180/9 c=20

OpenStudy (anonymous):

what is the answer?????????

OpenStudy (bakonloverk):

20

OpenStudy (anonymous):

thanks

OpenStudy (bakonloverk):

ur welcome